OpenStudy (anonymous):

Calculate the area of a circle with a radius of 2 inches and a circle with a radius of 4 inches. Leave your answers in terms of pi. When the radius doubles does the area of the circle double

OpenStudy (anonymous):

A 1= pi ( 2 )^2 A1 = 4 pi A2= 16 pi No

OpenStudy (anonymous):

no , it quadruples: A=pi*r^2 A1=4pi A2=16pi proof that it quadruples: A1=pi*r^2 A2=pi*(2r)^2 A2=pi*4r^2